3. A gas expands in volume from 26.7 mL to 89.3 mL at constant temperature. Calculate the work done (in joules) if the gas expands.

  a.) against a vacuum

  b.) against a constant pressure if 1.5 atm and,

  c.) against a constant pressure of 2.8 atm


1
Expert's answer
2022-03-19T02:44:18-0400

W = P∆V = P(V2-V1)

a) because the system is in a vacuum, external pressure is zero

=0(89.3-26.7) =0J

b) P= 1.5atm

= 1.5(89.3-26.7) = 93.9J

c) P = 2.8atm

= 2.8(89.3-26.7) = 175.28